OpenStudy (anonymous):

Help please I will medal and fan! A rectangle has an area of 36 square centimeters and a length of 9 centimeters. What is the width of the rectangle?

OpenStudy (anonymous):

4 centimeters

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

a=wl w=a/l So... 36 = 9 * w 4 = w Divide both sides by 9

OpenStudy (anonymous):

So it's 4

OpenStudy (anonymous):

Yes. Because 36/9 is 4

OpenStudy (anonymous):

Ok. thank you. Can you help me with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What is the area of the trapezoid? A. 320 square feet B. 160 square feet C. 72 square feet D. 56 square feet

OpenStudy (anonymous):